OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (1, 9); Slope: 5 A. y – 1 = 5(x + 9) B. y – 9 = 5(x + 1) C. y – 9 = 5(x – 1) D. y + 9 = 5(x – 1)

OpenStudy (anonymous):

If you help me then I will fan you and give you a medal.

OpenStudy (zpupster):

Point-Slope Equation of a Line: y-y1= m(x-x1) m=slope you get x1 and y1 from the given coordinates (x1,y1) plug your values in the equation what answer did you get?

TheSmartOne (thesmartone):

\[y-y _{1}=m(x-x _{1})\]

TheSmartOne (thesmartone):

x1 is 1 y1 is 9 m is 5

TheSmartOne (thesmartone):

\[y-9=5(x-1)\]

TheSmartOne (thesmartone):

Which should be C.

TheSmartOne (thesmartone):

@doresmith

OpenStudy (anonymous):

Thank you. Can you help me with a few more please?

TheSmartOne (thesmartone):

Post them in anohter question.

OpenStudy (anonymous):

okay.

TheSmartOne (thesmartone):

Don't forget to give the person who helped you most the medal.